two banquet halls are considered for a wedding reception. hall A charges s fixed cost of $6000 plus $125 per guest. hall B charges a fixed cost of $7500 plus $100 per guest. using the system of linear equations, find out when the two halls are the same cost
Answer by Boreal(15235)

guests=x
cost for A=125x+6000
cost for B=100x+7500
When the costs are equal, the equations are equal
125x+6000=100x+7500
25x=1500
x=60 guests--at that number the costs will be identical, and they will be $13,500.
A will be 125(60)+6000=$13,500
B will be 100(60)+7500=$13,500